# Connect the Dots: Knowledge Graph–Guided Crawler Attack on Retrieval-Augmented Generation Systems

### v1

@@ 6.2 Effectiveness (RQ1)
Dataset Method SuccessRate Similarity Rouge-L TREC-COVID RAGTheif 0.1129 0.4920 0.1547 IKEA 0.2247 0.4779 0.1730 RAGCrawler 0.3839 0.6098 0.2408 SciDocs RAGTheif 0.0486 0.4275 0.1131 IKEA 0.0179 0.4829 0.1277 RAGCrawler 0.3810 0.5900 0.2285 NFCorpus RAGTheif 0.0447 0.5156 0.1063 IKEA 0.4215 0.6064 0.2013 RAGCrawler 0.5259 0.6992 0.2334
Reconstruction Fidelity. To assess the practical utility of the extracted knowledge, we measure its effectiveness in a downstream task: building a surrogate RAG system. A surrogate RAG constructed from the knowledge base extracted by RAGCrawler demonstrates markedly superior performance. As reported in Tab. II, it achieves answer success rates ranging from 38.1% to 52.6% on official query sets. Moreover, it obtains the highest scores in embedding similarity (up to 0.699) and ROUGE-L (up to 0.240), confirming the high quality and functional value of the recovered knowledge. This operational effectiveness stems from the comprehensive and semantically integral knowledge base that RAGCrawler extracts. In contrast, surrogate systems built from the knowledge extracted by IKEA and RAGThief perform poorly. This highlights that their extracted content is not only less complete but also less useful for reconstructing the original system’s capabilities.
Takeaway. RAGCrawler outperforms prior methods across all datasets and generators, achieving an average coverage rate of 66.8%, compared to 46.1% for IKEA and 9.6% for RAGThief. It also obtains a superior average semantic fidelity of 0.605. Furthermore, surrogate RAG systems built from its extractions yield the highest answer success rates, confirming that RAGCrawler constitutes a more potent and realistic threat to RAG systems.
@@ 6.3 Retriever Sensitivity (RQ2)
RAGCrawler demonstrates robustness and adaptability, maintaining its superior performance. As shown in Table III for Llama-3-8B generator, RAGCrawler achieves an average coverage rate (CR) of 76.6% and a semantic fidelity (SF) of 0.596. Notably, its coverage improves compared to its performance with the BGE retriever (69.0%, from Table I), showcasing its ability to effectively capitalize on the document set surfaced by any underlying retriever.
TABLE III: Coverage rate (CR) and semantic fidelity (SF) of attacks on RAG using GTE as retriever (Llama-3-8B generator). RAGCrawler remains dominant.
Dataset Metric RAGTheif IKEA RAGCrawler TREC-COVID CR 0.191 0.391 0.765 SF 0.453 0.565 0.610 SciDocs CR 0.162 0.368 0.833 SF 0.345 0.495 0.539 NFCorpus CR 0.386 0.622 0.738 SF 0.589 0.648 0.671 Healthcare CR 0.388 0.671 0.728 SF 0.548 0.576 0.567
In contrast, while the baseline methods also see a performance uplift with the GTE retriever (IKEA’s average coverage increases to 53.9% and RAGThief’s to 28.2%), they still remain far less effective. This indicates that while a different retriever may surface “easier” content for all methods, the heuristic-driven approaches of the baselines are fundamentally less capable of exploiting this opportunity to the same degree as RAGCrawler’s strategic exploration.
Takeaway. RAGCrawler ’s effectiveness is not contingent on a specific retriever. When the victim RAG employs a GTE retriever, RAGCrawler achieves an average coverage of 76.6%, substantially outperforming both IKEA (53.9%) and RAGThief (28.2%). Its strategic, graph-based exploration makes it a more versatile and potent threat across diverse RAG system implementations.
@@ 6.4 Agent Sensitivity (RQ3)
RAGCrawler demonstrates high effectiveness irrespective of the agent’s scale. Even when powered by the smaller Qwen-2.5-7B-Instruct agent, RAGCrawler continues to achieve the highest performance across all metrics, as shown in Table IV. It reaches an average coverage of 68.9% and a semantic fidelity of 0.596, results that are comparable to, and in some cases even exceed, its performance with the larger agent (e.g., 54.2% coverage on TREC-COVID with Qwen vs. 49.4% with Doubao). This resilience stems from our framework’s design, which decouples high-level strategic planning from low-level content generation. The knowledge graph and strategy scheduler dictate the overall attack trajectory, requiring the LLM agent merely to execute localized and well-defined sub-tasks.
TABLE IV: Coverage rate (CR) and semantic fidelity (SF) of each attacks when using Qwen-2.5-7B-Instruct as attacker’s agent (victim RAG: Llama-3-8B generator, BGE retriever). Best results are in bold. RAGCrawler continues to attain the highest coverage and fidelity under a smaller model.
Dataset Metric RAGTheif IKEA RAGCrawler TREC-COVID CR 0.271 0.183 0.542 SF 0.453 0.499 0.610 SciDocs CR 0.314 0.559 0.675 SF 0.345 0.495 0.539 NFCorpus CR 0.386 0.622 0.738 SF 0.589 0.648 0.671 Healthcare CR 0.414 0.687 0.799 SF 0.548 0.576 0.567
In comparison, the baseline methods remain significantly less effective. With the Qwen-2.5-7B-Instruct agent, IKEA achieves an average coverage of 51.3%, while RAGThief reaches 34.6%. Although RAGThief sees a slight performance improvement (likely because the constrained divergent ability of smaller model make it less prone to the off-topic drift), it still lags far behind RAGCrawler.
Takeaway. The effectiveness of RAGCrawler is largely independent of the LLM agent. When using a smaller model such as Qwen-2.5-7B-Instruct, it achieves an average coverage of 68.9%, decisively outperforming baselines. This highlights that the primary strength of our attack lies in its strategic planning, not the raw capability of agent model.
@@ 6.5 Defense Robustness (RQ4)
TABLE V: Coverage rate (CR) and semantic fidelity (SF) of attacks under query rewriting (victim RAG: Llama-3-8B generator, BGE retriever). Best results are in bold. RAGCrawler still achieves the highest coverage and fidelity in this setting, indicating resilience to query rewriting.
Dataset Metric RAGTheif IKEA RAGCrawler TREC-COVID CR 0.381 0.241 0.601 SF 0.519 0.537 0.591 SciDocs CR 0.561 0.542 0.743 SF 0.442 0.479 0.507 NFCorpus CR 0.664 0.489 0.854 SF 0.633 0.618 0.687 Healthcare CR 0.709 0.595 0.767 SF 0.556 0.564 0.577
Query Rewriting. In this configuration, the victim RAG system employs an LLM to rewrite incoming queries, aiming to clarify user intent and neutralize potential adversarial patterns before retrieval and generation. RAGCrawler exhibits exceptional resilience against this defense, maintaining superior performance metrics. As detailed in Table V, RAGCrawler achieves an average coverage rate of 74.1% and semantic fidelity of 0.591. Although intended as a safeguard, the query rewriting process is paradoxically exploited by our method to enhance extraction. By explicitly refining the query’s semantic intent, the rewriter enables the retriever to surface a more relevant and diverse set of documents than the original input would yield. RAGCrawler’s adaptive planner capitalizes on this context to accelerate corpus exploration; for instance, on the NFCorpus dataset, coverage increases from 73.8% to 85.4% when this defense is active.
@@
TABLE VI: Coverage rate (CR) and semantic fidelity (SF) of attacks under multi-query retrieval (victim RAG: Llama-3-8B generator, BGE retriever). Best results are in bold. RAGCrawler still achieves the highest coverage and fidelity, indicating resilience to multi-query retrieval.
Dataset Metric RAGTheif IKEA RAGCrawler TREC-COVID CR 0.326 0.189 0.474 SF 0.525 0.523 0.581 SciDocs CR 0.352 0.508 0.697 SF 0.364 0.494 0.514 NFCorpus CR 0.392 0.540 0.849 SF 0.588 0.631 0.692 Healthcare CR 0.562 0.605 0.761 SF 0.569 0.580 0.587
Multi-query Retrieval. We next evaluate against multi-query retrieval, which expands each query into several variants whose retrieved results are re-ranked and fused, a process that can influence attack surface. Once again, RAGCrawler excels, achieving the highest average coverage of 69.5% and a semantic fidelity of 0.593 (Table VI). This mechanism provides the attacker with a more diverse set of retrieved documents from different semantic clusters. The global graph in RAGCrawler is uniquely positioned to exploit this; it integrates this diverse information to build a more comprehensive map of the corpus, thereby generating more effective and far-reaching follow-up queries.
The baseline methods, however, are not equipped to fully leverage this enriched context. IKEA and RAGThief attain coverage rates of 46.1% and 40.8%, respectively. Their localized strategies are unable to fully synthesize the information from the multiple retrieval results.
Takeaway. RAGCrawler remains highly effective against practical RAG defenses. The results demonstrate that common defenses designed to sanitize inputs or strenghthen retrieval can be subverted by a strategic attacker and may even amplify the extraction threat, highlighting the need for more advanced safeguards.
@@
6.6 Ablation Study (RQ5)
We first study the influence of score cache and victim retrieval depth. More ablations on modules and hyperparameter choices are provided in Appendix 9.3.
@@ 6.6 Ablation Study (RQ5)
Score Cache. The Strategy Scheduler module includes a score cache to avoid redundant computations of the CMG for similar states. We measured the impact of this optimization on computational overhead. As shown in Fig. 9, across the evaluated datasets the attacker would need to update over 1.67 million scoring operations without caching, versus only about 0.63 million with our cache (a 62% reduction in workload). In practice, this translates to a faster attack runtime and lower cost, enabling the attacker to scale to larger corpora or tighter time budgets. The benefit of caching grows with the number of queries, since later attack rounds re-use many of the same entity-relation state components.
Victim Retrieval Depth. We also analyzed sensitivity of RAGCrawler to the victim RAG’s retrieval depth (k), the number of documents retrieved per query. As illustrated in Fig. 10 for the TREC-COVID dataset, a larger $k$ allows the attacker to acquire knowledge more rapidly; for instance, coverage within 1,000 queries increases from approximately 28% at k=5 to over 60% at k=20. However, this effect is subject to diminishing returns. As $k$ grows, the additional documents retrieved are increasingly likely to have semantic overlap with previously seen content, providing less novel information for the planner to exploit.
@@ 7 Discussion and Related Work
Cost Analysis. RAGCrawler is highly cost-effective, with a monetary cost of only \$0.33–\$0.53 per dataset (Tab. VIII) and is reducible to near-zero using open-source models (Sec. 6.4). This minimal expenditure creates a stark economic asymmetry against the high development costs of victim RAG systems [adasci2024ragcost]. This asymmetry poses a dual threat: enabling low-cost service piracy by reconstructing the knowledge base (Tab. II) and exposing victims to severe legal and privacy liabilities under regulations like HIPAA and GDPR if sensitive data is exfiltrated [annas2003hipaa, voigt2017eu].
Defenses. Our attack highlights the inherent limitations of defenses that rely on analyzing queries in isolation. Its use of individually benign queries makes it difficult to counter with conventional tools like guardrails or query rewriting without significant performance penalties. The core challenge is that the malicious pattern is an emergent property of the entire interaction sequence, not an attribute of any single query. This exposes a critical blind spot in static security models and argues for a pivot towards dynamic, behavior-aware defenses. Based on this analysis, we believe query provenance analysis [li2025query, han2020unicorn, li2023nodlink, milajerdi2019holmes] holds significant promise. Because this technique is designed to track relationships across entire sequences, it is theoretically capable of identifying the progressive, goal-oriented patterns of an extraction attack. Verifying its practical effectiveness would be a valuable next step for the research community.
Security Risks in RAG. This paper concentrates on Data Extraction Attacks against RAG systems, where an adversary queries the system to steal or reconstruct its private knowledge base [qifollow, jiang2024rag, wang2025silent, di2024pirates, cohen2024unleashing]. RAG systems are also susceptible to other security risks [lasso2024ragsecurity, ni2025towards2]. Membership Inference Attacks (MIA) enable an adversary to determine if a specific document is present in the private corpus, thereby exposing the existence of sensitive records [liu2025mask, naseh2025riddle, anderson2025my, feng2025ragleak]. Furthermore, in Corpus Poisoning attacks, an adversary injects malicious data into the RAG knowledge base [zou2025poisonedrag, chaudhari2024phantom, cho2024typos, shafran2024machine]. Retrieval of this corrupted data can manipulate the system’s output, allowing the attacker to propagate misinformation or harmful content.

### v2

@@ 7.2 Effectiveness (RQ1)
Dataset RAGThief IKEA RAGCrawler TREC-COVID >5,000>5{,}000 (29.8%) >5,000>5{,}000 (19.6%) 4,040 SciDocs >5,000>5{,}000 (7.2%) >5,000>5{,}000 (56.0%) 1,163 NFCorpus >5,000>5{,}000 (16.5%) >5,000>5{,}000 (64.0%) 595 Healthcare >5,000>5{,}000 (53.0%) 1,231 568
[…]
The full sweep in Tab. 10 highlights the gap. While IKEA can be competitive at low targets on easier corpora (e.g., Healthcare $Q_{0.5}=212$ vs. 209), it saturates as $\gamma$ increases and fails to reach 90% on any dataset within $Q_{\max}$. In contrast, RAGCrawler reaches 90% on SciDocs, NFCorpus, and Healthcare with 3,330, 1,753, and 2,434 queries, respectively, and attains 72.5% final coverage on TREC-COVID, more than doubling the best baseline’s (29.8% for RAGThief).
Table 3: Reconstruction fidelity of surrogate RAG systems built from stolen knowledge-base content (victim: Llama-3-8B generator, BGE retriever). We report success rate (non-refusal), answer embedding similarity, and ROUGE-L against victim outputs. Best results are in bold. Surrogates built from RAGCrawler’s stolen corpus match the victim best.
Dataset Method Success Rate Answer Sim. ROUGE-L TREC-COVID RAGThief 0.1129 0.4920 0.1547 IKEA 0.2247 0.4779 0.1730 RAGCrawler 0.3839 0.6098 0.2408 SciDocs RAGThief 0.0486 0.4275 0.1131 IKEA 0.0179 0.4829 0.1277 RAGCrawler 0.3810 0.5900 0.2285 NFCorpus RAGThief 0.0447 0.5156 0.1063 IKEA 0.4215 0.6064 0.2013 RAGCrawler 0.5259 0.6992 0.2334
Reconstruction Fidelity. To test whether the stolen corpus enables service replication, we build surrogate RAG systems from each method’s stolen outputs and evaluate them on held-out query sets. Table 3 shows that surrogates built from RAGCrawler achieve substantially higher success rates (38.1%-52.6%) and consistently higher answer similarity and ROUGE-L, with peaks of 0.6992 similarity and 0.2408 ROUGE-L. Notably, on SciDocs the IKEA-based surrogate has a very low success rate (0.0179) despite IKEA reaching 51.3% coverage under the 1,000-query budget in Table 1, indicating that partial theft concentrated in a limited region does not necessarily translate into functional replication. Overall, the reconstruction results align with the coverage and fidelity trends: broader exploration coupled with higher semantic alignment yields stolen corpora that are materially more useful for building a substitute service.
Takeaway. Across datasets and generators, RAGCrawler delivers the strongest stealing breadth, semantic fidelity, and query efficiency, and the resulting stolen content transfers into markedly higher-fidelity surrogate reconstruction.
@@ 7.3 Retriever Sensitivity (RQ2)
Swapping the retriever shifts the attack surface for all methods, but RAGCrawler remains the most effective extractor in terms of coverage. With GTE, RAGCrawler achieves 76.6% average coverage across datasets, compared with 53.9% for IKEA and 28.2% for RAGThief (Table 4). The effect of the retriever is not uniform across corpora: relative to BGE, RAGCrawler gains 27.1 points on TREC-COVID (49.4% to 76.5%) and 17.2 points on SciDocs (66.1% to 83.3%), but drops by 5.9 and 7.9 points on NFCorpus and Healthcare, respectively. This variation suggests that retriever choice can materially change which parts of a corpus are easy to surface, yet RAGCrawler consistently maintains high coverage (at least 72.8%) across all four corpora under GTE.
Table 4: Coverage rate (CR) and semantic fidelity (SF) of attacks when the victim uses GTE as retriever (generator: Llama-3-8B; 1,000-query budget). Best results are in bold. RAGCrawler remains dominant.
Dataset CR SF RAGThief IKEA RAGCrawler RAGThief IKEA RAGCrawler TREC-COVID 0.191 0.391 0.765 0.453 0.565 0.610 SciDocs 0.162 0.472 0.833 0.345 0.495 0.539 NFCorpus 0.386 0.622 0.738 0.589 0.648 0.671 Healthcare 0.388 0.671 0.728 0.548 0.576 0.567 Average 0.282 0.539 0.766 0.484 0.571 0.597
Semantic fidelity remains broadly stable. Averaged over datasets, RAGCrawler attains SF 0.597, exceeding IKEA (0.571) and RAGThief (0.484). The only exception is Healthcare, where IKEA achieves marginally higher SF (0.576 vs. 0.567), indicating that retriever changes can also alter per-document alignment quality in specific domains.
Takeaway. RAGCrawler generalizes across retrievers: changing the victim retriever can affect absolute difficulty and the coverage–fidelity balance, but RAGCrawler remains the strongest overall threat, especially in coverage.
Table 5: Coverage rate (CR) and semantic fidelity (SF) of attacks when using Qwen-2.5-7B-Instruct as the attacker-side agent (victim: Llama-3-8B generator, BGE retriever; 1,000-query budget). Best results are in bold. RAGCrawler remains dominant.
Dataset CR SF RAGThief IKEA RAGCrawler RAGThief IKEA RAGCrawler TREC-COVID 0.271 0.183 0.542 0.469 0.499 0.570 SciDocs 0.099 0.489 0.675 0.298 0.487 0.507 NFCorpus 0.314 0.559 0.834 0.559 0.646 0.678 Healthcare 0.414 0.687 0.799 0.549 0.576 0.584 Average 0.275 0.480 0.713 0.469 0.552 0.585
@@ 7.4 Agent Sensitivity (RQ3)
RAGCrawler remains effective with a smaller attacker model. With Qwen-2.5-7B-Instruct (victim: Llama-3-8B, BGE), RAGCrawler achieves 71.3% average coverage with 0.585 semantic fidelity (Table 5). Compared to the Doubao agent under the same victim setting, the per-dataset coverage changes are small (within a few percentage points), suggesting that the global scheduler and graph state dominate the attack trajectory, while the LLM agent primarily executes localized, well-specified generation tasks.
In comparison, the baseline methods remain significantly less effective. With the Qwen-2.5-7B-Instruct agent, IKEA achieves an average coverage of 48.0%, while RAGThief reaches 27.5%. Although RAGThief sees a slight performance improvement (likely because the constrained divergent ability of smaller model make it less prone to the off-topic drift), it still lags far behind RAGCrawler.
Takeaway. Attack effectiveness for RAGCrawler is largely agent-agnostic, reinforcing that its advantage comes from planning and redundancy control rather than relying on a high-capability attacker LLM.
Table 6: Coverage rate (CR) and semantic fidelity (SF) under practical RAG variants (victim: Llama-3-8B generator, BGE retriever, 1,000-query budget). Best results are in bold. RAGCrawler remains the strongest attack.
Query Rewriting Dataset CR SF RAGThief IKEA RAGCrawler RAGThief IKEA RAGCrawler TREC-COVID 0.381 0.241 0.601 0.519 0.537 0.591 SciDocs 0.561 0.542 0.743 0.442 0.479 0.507 NFCorpus 0.664 0.489 0.854 0.633 0.618 0.687 Healthcare 0.709 0.595 0.767 0.556 0.564 0.577 Average 0.579 0.467 0.741 0.538 0.550 0.591 Multi-query Retrieval Dataset CR SF RAGThief IKEA RAGCrawler RAGThief IKEA RAGCrawler TREC-COVID 0.326 0.189 0.474 0.525 0.523 0.581 SciDocs 0.352 0.508 0.697 0.364 0.494 0.514 NFCorpus 0.392 0.540 0.849 0.588 0.631 0.692 Healthcare 0.562 0.605 0.761 0.569 0.580 0.587 Average 0.408 0.461 0.695 0.512 0.557 0.594
@@ 7.5 Defense Robustness (RQ4)
Query Rewriting. In this configuration, the victim RAG system employs an LLM to rewrite incoming queries, aiming to clarify user intent and neutralize potential adversarial patterns before retrieval and generation. RAGCrawler exhibits exceptional resilience against this defense, maintaining superior performance metrics. As detailed in Table 6, RAGCrawler achieves an average coverage rate of 74.1% and semantic fidelity of 0.591. Although intended as a safeguard, the query rewriting process is paradoxically exploited by our method to enhance theft. By explicitly refining the query’s semantic intent, the rewriter enables the retriever to surface a more relevant and diverse set of documents than the original input would yield. RAGCrawler’s adaptive planner capitalizes on this context to accelerate corpus exploration; for instance, on the NFCorpus dataset, coverage increases from 79.7% to 85.4% when this defense is active.
@@
Multi-query Retrieval. We next evaluate against multi-query retrieval, which expands each query into several variants whose retrieved results are re-ranked and fused, a process that can influence attack surface. Once again, RAGCrawler excels, achieving the highest average coverage of 69.5% and a semantic fidelity of 0.593 (Table 6). This mechanism provides the attacker with a more diverse set of retrieved documents from different semantic clusters. The global graph in RAGCrawler is uniquely positioned to exploit this; it integrates this diverse information to build a more comprehensive map of the corpus, thereby generating more effective and far-reaching follow-up queries.
The baseline methods, however, are not equipped to fully leverage this enriched context. RAGThief and IKEA attain coverage rates of 40.8% and 46.1%, respectively. Their localized strategies are unable to fully synthesize the information from the multiple retrieval results.
Takeaway. RAGCrawler remains highly effective against practical RAG defenses. The results demonstrate that defenses designed to sanitize inputs or strengthen retrieval can be subverted by a strategic attacker and may even amplify the stealing threat, calling for more advanced safeguards.
7.6 Ablation Study (RQ5)
We analyze how attacker assumptions, efficiency techniques, and victim-side hyperparameters influence RAGCrawler’s effectiveness. More ablations on modules and hyperparameter choices are provided in Appendix E.2.
Table 7: Coverage rate (CR) with a provided topic prior vs. a no-prior attacker (victim: Llama-3-8B generator, BGE retriever). Best results are in bold. RAGCrawler remains best and shows small sensitivity to topic-prior removal.
TREC-COVID SciDocs NFCorpus Healthcare Method Prior No-prior Prior No-prior Prior No-prior Prior No-prior RAGThief 0.131 0.128 0.053 0.040 0.061 0.123 0.361 0.290 IKEA 0.161 0.391 0.513 0.108 0.503 0.536 0.687 0.705 RAGCrawler 0.494 0.426 0.661 0.613 0.797 0.825 0.807 0.755
Robustness to topic prior. Our main threat model assumes a coarse topic phrase [cohen2024unleashing, jiang2024rag, wang2025silent], which is typically exposed by production RAG services. We further evaluate a weaker attacker that starts with no topic prior and infers a coarse 3–8 word topic phrase using a single benign probe query (counted toward the budget). Table 7 shows that RAGCrawler remains the best-performing attack in this setting, with only modest CR changes relative to the topic-prior setting. Appendix D details the one-shot probing procedure and analyzes the quality and variability of the inferred topic phrases.
@@ 7.6 Ablation Study (RQ5)
[…]
@@
Victim Retrieval Depth. We vary the victim retrieval depth $k$ to study how much per-query context expansion affects attack performance. On TREC-COVID, increasing $k$ substantially accelerates early coverage growth: within 1,000 queries, coverage rises from roughly 28% at $k=5$ to above 60% at $k=20$ (Fig. 10). However, the curve exhibits diminishing returns as $k$ increases, consistent with additional retrieved documents being increasingly redundant with already exposed content. This result highlights a security implication of aggressive retrieval: larger $k$ increases the amount of unique corpus material that can be exposed per user query.
@@ 8 Discussion and Related Work
Cost Analysis. We separate (i) attacker-side orchestration cost (tokens for running RAGCrawler) from (ii) victim-side service cost (per-query fees). Tab. 9 shows RAGCrawler incurs only \$0.33–\$0.53 per dataset under Doubao-1.5-lite-32K pricing, and becomes near-zero with open-source models (Sec. 7.4) aside from GPU hosting. Victim-side fees scale linearly with issued queries (e.g., $Q_{0.7}$ in Tab. 2). The low attacker overhead creates strong economic asymmetry against RAG development cost [adasci2024ragcost], enabling low-cost service piracy.
Scalability. Although production RAG corpora may contain millions of documents, RAGCrawler scales because each turn extracts a compact typed delta from the single observed answer to update the attacker-side KG, keeping per-turn LLM cost independent of $|\mathcal{D}|$. The main bottleneck is accumulated-state maintenance (semantic merging and score updates), which we mitigate with caching and lazy recomputation (62% fewer score updates; Fig. 9). For larger budgets, ANN-backed deduplication [malkov2018efficient, yang2025effective] and out-of-core graph and score storage [ai2017squeezing, zhang2018wonderland, xu2020hybrid] offer orthogonal engineering extensions.
Defenses. Our attack highlights the inherent limitations of defenses that rely on analyzing queries in isolation (e.g., guardrails or query rewriting). Moreover, when we evaluate an LLM-based intent detector, it flags only 1.07% of the attack queries as suspicious. The core challenge is that the malicious pattern is an emergent property of the entire interaction sequence, not an attribute of any single query. This exposes a critical blind spot in static security models and argues for a pivot towards dynamic, behavior-aware defenses. Based on this analysis, we believe query provenance analysis [li2025query, han2020unicorn, li2023nodlink, milajerdi2019holmes] holds significant promise. Verifying its practical effectiveness would be a valuable next step for the research community.
Security Risks in RAG. This paper concentrates on knowledge-base stealing attacks against RAG systems [qifollow, jiang2024rag, wang2025silent, di2024pirates, cohen2024unleashing]. RAG systems are also susceptible to other security risks [lasso2024ragsecurity, ni2025towards2]. Membership Inference Attacks (MIA) enable an adversary to determine if a specific document is present in the corpus, thereby exposing the existence of sensitive records [liu2025mask, naseh2025riddle, anderson2025my, feng2025ragleak]. Furthermore, in Corpus Poisoning attacks, an adversary injects malicious data into the RAG knowledge base [zou2025poisonedrag, chaudhari2024phantom, cho2024typos, shafran2024machine, zhong2023poisoning, liu2025poisoned, ben2024gasliteing, liang2025graphrag]. Retrieval of this corrupted data can manipulate the system’s output, allowing the attacker to propagate misinformation or harmful content.